# MobileVLM: A Vision-Language Model for Better Intra- and Inter-UI Understanding

### v1

@@ Abstract
Recently, mobile AI agents based on VLMs have gained increasing attention. These works typically utilize VLM pre-trained on general-domain data as a foundation, fine-tuning it on instruction-based mobile datasets. However, the proportion of mobile UI in general pre-training data is very low. Moreover, the general pre-training task does not particularly consider the characteristics of mobile UI. Therefore, directly applying such pre-trained models for mobile UI instruction fine-tuning will not yield the desired performance. In this paper, we propose MobileVLM for Chinese UI manipulation. On top of the general pre-training model, two additional pre-training stages are implemented with four specific tasks to enhance both intra-and inter-UI understanding. In addition, a large Chinese mobile UI corpus, named Mobile3M, is built from scratch to compensate for the lack of relevant data. Besides 3 million static UI pages, it also contains directed graph structures formed by real-world UI transition actions. Experimental results show MobileVLM excels on both in-house test sets and public mobile benchmarks, outperforming existing VLMs.
@@ 1 Introduction
Mind2Web Deng et al. (2024) English Web 2,350 7.3 HTML Chain n/a
@@ 2.1 Mobile UI Dataset
Table 1 provides a comparison of multiple mobile UI datasets. At the top of the table are several "point" datasets. Each data instance in these datasets contains only one page, along with different fine-grained tasks and corresponding answers. Rico Deka et al. (2017) is a large-scale Android UI dataset and has been widely used as a primary data source for UI modeling research. UIBert Bai et al. (2021) release two new datasets extended from Rico. Ferret-UI You et al. (2024) uses the UI detection model Zhang et al. (2021b) to annotate fine-grained elements in Android and iPhone screens. However, these datasets only focus on the elements and layout within a single page, so it is difficult for them to capture the complete process of users using the app.
To better reflect user behavior, several "chained" mobile UI data sets have been released Sun et al. (2022); Deng et al. (2024); Burns et al. (2021), as shown at the bottom of Table 1. Each "chain" of data in these datasets consists of a sequence of action-UI pages. The UI page includes a screenshot and a structured document. AITW Rawles et al. (2023) is one of the largest UI control datasets with 5 subsets and 715K episodes. Auto-UI Zhan and Zhang (2023) further filters the GoogleApps subset in the AITW dataset, leaving 152K episodes. However, these datasets only provide pages and OCR text, missing structural documents, which makes it difficult for the VLM model to learn the ability to align image and text modalities. As shown in Figure 2, compared with chained datasets, Mobile3M’s graph structure can better capture the relationship between different pages in the app.
@@ 3.1 Background
UI Page: We selected 49 Chinese apps from the App Storehttps://sj.qq.com/, ensuring that each app had at least 10 million downloads. The apps are installed and run on the emulator, and we use Appiumhttps://appium.io/docs/en/latest/ to collect UI pages. The UI page includes a screenshot and an XML document. The XML document describes the structure and content of a UI page, including elements like buttons and text boxes, as well as layout information such as bounding boxes and hierarchical trees. Figure 11 in the Appendix shows an example of a screenshot and an XML document for the app "QQMusic". The XML document can be parsed to produce a list of elements. As shown in the task (c) of Figure 3, each element contains a name and a bounding box.
@@ 3.2 Data Collection
Inspired by APPAgent Yang et al. (2023), we use the random walk algorithm to explore apps. The algorithm iteratively interacts with every element on each UI page and records the page transition states. The exploration results for each app can be represented as a directed graph, in which each edge represents an action and each node represents a UI page. The "action trace" of a UI page is defined as the shortest sequence of actions from the app’s homepage to that page. The ID of each action in the trace is combined to create a unique identifier for the page called “page name”. In Figure 2, the algorithm executes "click(Edit)" from the "Baicizhan0" page to enter the Edit page. The ID of this action is 1, so the name identifier of the Edit page is assigned as Baicizhan0_1.
Since the UI pages from real-world apps may change as the app evolves. During exploration, if the head node stored several days before is simply approximated to the current head node, UI changes may be mistaken for the edge’s action, causing data errors. Therefore, for each page, the algorithm will save screenshots and XML documents of each step in its entire "action trace". Taking "Baicizhan0_1_25" in Figure 2 as an example, this node contains 3 UI pages, 3 XML documents, and an "action trace" consisting of 2 actions.
We adopt the breadth-first method (BFS) to explore apps. Compared with the depth-first method (DFS), BFS better covers app functions and shortens the action sequence when exploring new nodes. As shown in Figure 2, the algorithm will first explore "Baicizhan0_10" instead of "Baicizhan0_1_25". The task-oriented exploration method Yang et al. (2023) heavily relies on the performance of VLM. However, current VLMs can be costly and may perform poorly with third-party apps. In addition, the task-oriented method may cause some infrequently used pages and app functions to be overlooked.
@@ 3.3 Method Optimization
To improve exploration efficiency, we propose a "unique page" mechanism. Every time a new page is explored, we use BM25 Robertson et al. (2009) to retrieve the top 5 nodes in the current app graph that are closest to the XML document of the page. The algorithm compares the new page with each of these five pages to determine if they are similar pages. The threshold of the similar coefficient is Element Diff<5 & Pixel Diff <30%. Here, Element Diff is the number of different elements between two UI pages and Pixel Diff is the pixel difference between two screenshots. If no similar page is found in the current graph, the new page will be treated as a unique page and added to the app graph. As shown in Figure 2, click the "Back" button on "Baicizhan0_1_24", and the generated "Baicizhan0_1_24_113" and "Baicizhan0" are equivalent pages. We add a directed edge from the previous page "Baicizhan0_1_24" to the similar page "Baicizhan0" in the graph. The algorithm will not treat the "Baicizhan0_1_24_113" as a new explorable node.
The benefits of this mechanism are threefold: 1. This greatly reduces the exploration space of each app. Taking "ctrip" as an example, our exploration process produced 187,079 UI pages with an average steps of 6.5. Without the "unique page" mechanism, pages of this magnitude cannot even cover all possibilities of 4-step exploration. 2. This converts the tree structure exploration results into a graph structure. Different pages can reach "Baicizhan0_1" by clicking "Edit", "OK", and scrolling. This helps the agent learn the functions of different UI elements. 3. This helps prevent the occurrence of cyclic action sequences. The "unique page" mechanism can detect and prune them.
To balance the distribution of different actions in the dataset, during random walks, we give priority to the input action. We provide 10 related keywords for each app. When executing the input action, the algorithm can randomly select a keyword to input. For scroll actions, the algorithm can choose a direction to scroll from "up, down, left, and right".
@@ 3.4 Dataset Statistics
Among the 49 selected apps, we ensure that each main category in AppStore contains at least 2 apps. Figure 4 shows the data distribution of the Mobile3M dataset. The most common application categories in the dataset are "Travel", "Living" and "Shopping". As shown in the figure, Mobile3M covers multiple categories, and the amount of data in each category is relatively balanced, which ensures that the dataset is versatile and diverse.
@@ 4.3 Model Architecture
Table 4: Main Result(%). Prior published best results are marked with an underline. The BC results are from (Sun et al., 2022), results with * are from Ding (2024), and results with # stem from (Zhan and Zhang, 2023). Considering the testing costs, Qwen-VL-Max and GPT-4o were conducted on a random sample of 10% of Auto-UI.
@@ A.3 Few-shot Prompt
For the Grounding Task, we use this few-shot prompt to guide the LLMs in answering the questions:
Here are three examples and a question. You need to help me find the location of the text I’m looking for in the image and output its bounding box. Please note that the coordinates are relative to the top-left corner of the image. Here are the three examples:
@@
For the Navigation Task, we use this few-shot prompt to guide the LLMs in answering the questions:
Here are two examples and a question. You need to tell me which control to interact with to navigate from the first image to the second image. There are three actions: click, input, and scroll. Here are the two examples:
@@
For the OCR Task, we use this few-shot prompt to guide the LLMs in answering the questions:
@@
    Question: Image: <img>/home/corpus/test_515/few_shot/ocr/ctrip0_1_36_313_3617_1988_8364 _4440_1566/ctrip0_1_36_313_3617_1988_8364_4440_1566-screen.png</img> Explain how the elements on the current page support its content and functionality.
@@
    Question: Image: <img>/home/corpus/test_515/few_shot/ocr/baicizhan0_1_24_113_159_165_156 _206/baicizhan0_1_24_113_159_165_156_206-screen.png</img> Summarize the layout and interaction methods of each part of the page.
@@
    <ref>Added</ref><box>[613,524][682,558]</box>
@@ B.1 Background
B.2 Graph Structure
@@
B.3 Category Details
Figure 7 shows a two-layered ring chart. The inner ring represents the data proportion by category, while the outer ring indicates the specific quantity proportion of each app. From this figure, it can be seen that, except for a few apps, most categories have a uniform distribution. For detailed quantity statistics, refer to Table 2.
Category APP Unique Nodes All Nodes Action Steps All Nodes (%)
Living anjuke 57,286 190,102 1,334,428 6.13%
Living wuba 38,667 147,009 903,586 4.74%
Living smarthome 12,595 42,961 304,816 1.39%
Travel ctrip 63,449 187,079 1,217,304 6.04%
Travel Qunar 42,462 161,005 1,211,015 5.20%
Shopping vipshop 72,468 168,531 1,036,086 5.44%
Shopping xiaomiShop 21,666 99,770 755,718 3.22%
Shopping duapp 18,925 38,926 223,379 1.26%
Transport didi 12,786 84,865 637,400 2.74%
Transport cainiao 20,593 73,132 480,223 2.36%
Transport gaodeMap 13,674 59,142 319,377 1.91%
Transport BaiduMap 13,552 54,322 280,498 1.75%
Browser UCMobile 40,618 88,220 615,049 2.85%
Browser baiduBrowser 36,016 70,282 401,348 2.27%
Browser QQBrowser 18,500 44,006 218,828 1.42%
Browser tencentnews 23,408 38,241 224,804 1.23%
System taptap 24,759 105,461 624,941 3.40%
System qqpimsecure 8,997 42,691 379,926 1.38%
System ludashi 2,773 32,474 219,804 1.05%
System qqdownloader 10,517 28,502 151,824 0.92%
System calculator 4,265 15,819 97,005 0.51%
System supercaculator 690 1,369 5,444 0.04%
Music ximalaya 34,995 103,395 577,032 3.34%
Music kugou 40,043 94,271 504,368 3.04%
Music QQmusic 5,545 17,539 64,211 0.57%
Reader seekbooks 15,902 70,266 563,882 2.27%
Reader QQReader 22,588 63,458 472,509 2.05%
Reader zhuishushenqi 14,737 63,210 392,903 2.04%
Reader pdfreader 495 1,507 5,211 0.05%
Social xiaohongshu 45,324 85,362 525,519 2.75%
Social zhihu 21,766 57,261 373,756 1.85%
Social QQ 7,051 20,600 141,969 0.66%
Education zuoyebang 19,884 70,661 507,146 2.28%
Education xiaoyuan 10,727 56,806 393,395 1.83%
Education youdao 8,756 35,121 206,035 1.13%
Education baicizhan 4,196 16,383 88,500 0.53%
Office wpsOffice 11,156 73,739 486,661 2.38%
Office netmail 5,544 32,308 260,682 1.04%
Office tonghuashun 6,410 30,722 163,297 0.99%
Office QQmail 712 1,590 4,597 0.05%
Video bili 46,080 91,891 471,940 2.97%
Video qqlive 12,497 22,601 99,677 0.73%
Video kuaishou 7,126 12,115 59,373 0.39%
Picture androidesk 28,432 59,228 418,773 1.91%
Picture mtxx 19,718 55,324 419,055 1.79%
Health medicinehelper 15,046 83,832 547,880 2.71%
Health keep 7,730 22,500 117,124 0.73%
Weather pureweather 25,252 79,283 610,695 2.56%
Weather cloudweather 1,956 3,904 19,339 0.13%
15 49 998,334 3,098,786 20,138,332 100%
Table 7: The data distribution in Mobile3M.

### v2

@@ Abstract
Recently, mobile AI agents based on VLMs have gained increasing attention. These works typically utilize VLM pre-trained on general-domain data as a foundation, fine-tuning it on instruction-based mobile datasets. However, the proportion of mobile UI in general pre-training data is very low. Moreover, the general pre-training task does not particularly consider the characteristics of mobile UI. Therefore, directly applying such pre-trained models for mobile UI instruction fine-tuning will not yield the desired performance. In this paper, we propose MobileVLM for Chinese UI manipulation. On top of the general pre-training model, two additional pre-training stages are implemented with four specific tasks to enhance both intra-and inter-UI understanding. In addition, a large Chinese mobile UI corpus, named Mobile3M, is built from scratch to compensate for the lack of relevant data. Besides 3 million static UI pages, it also contains directed graph structures formed by real-world UI transition actions. Experimental results show MobileVLM excels on both in-house test sets and public mobile benchmarks, outperforming existing VLMs. Dataset and Code are available at https://github.com/XiaoMi/mobilevlm.
@@ 1 Introduction
Mind2Web Deng et al. (2024b) English Web 2,350 7.3 HTML Chain n/a
@@ 2.1 Mobile UI Dataset
Table 1 provides a comparison of multiple mobile UI datasets. At the top of the table are several “point” datasets. Each data instance in these datasets contains only one page, along with different fine-grained tasks and corresponding answers. Rico Deka et al. (2017) is a large-scale Android UI dataset and has been widely used as a primary data source for UI modeling research. UIBert Bai et al. (2021) release two new datasets extended from Rico. Ferret-UI You et al. (2024) uses the UI detection model Zhang et al. (2021b) to annotate fine-grained elements in Android and iPhone screens. However, these datasets only focus on the elements and layout within a single page, so it is difficult for them to capture the complete process of users using the app.
To better reflect user behavior, several “chained” mobile UI data sets have been released Sun et al. (2022); Deng et al. (2024b); Burns et al. (2021); Deng et al. (2024a), as shown at the bottom of Table 1. Each “chain” of data in these datasets consists of a sequence of action-UI pages. The UI page includes a screenshot and a structured document. AITW Rawles et al. (2023) is one of the largest UI control datasets with 5 subsets and 715K episodes. Auto-UI Zhan and Zhang (2023) further filters the GoogleApps subset in the AITW dataset, leaving 152K episodes. However, these datasets only provide pages and OCR text, missing structural documents, which makes it difficult for the VLM model to learn the ability to align image and text modalities. As shown in Figure 2, compared with chained datasets, Mobile3M’s graph structure can better capture the relationship between different pages in the app.
@@ 3.1 Background
UI Page: We selected 49 Chinese apps from the App Storehttps://sj.qq.com/, ensuring that each app had at least 10 million downloads. The apps are installed and run on the emulator, and we use Appiumhttps://appium.io/docs/en/latest/ to collect UI pages. The UI page includes a screenshot and an XML document. The XML document describes the structure and content of a UI page, including elements like buttons and text boxes, as well as layout information such as bounding boxes and hierarchical trees. Figure 11 in the Appendix shows an example of a screenshot and an XML document for the app “QQMusic”. The XML document can be parsed to produce a list of elements. As shown in the task (c) of Figure 3, each element contains a name and a bounding box.
@@ 3.2 Data Collection
Inspired by APPAgent Yang et al. (2023), we use the random walk algorithm to explore apps. The algorithm iteratively interacts with every element on each UI page and records the page transition states. The exploration results for each app can be represented as a directed graph, in which each edge represents an action and each node represents a UI page. The “action trace” of a UI page is defined as the shortest sequence of actions from the app’s homepage to that page. The ID of each action in the trace is combined to create a unique identifier for the page called “page name”. In Figure 2, the algorithm executes “click(Edit)” from the “Baicizhan0” page to enter the Edit page. The ID of this action is 1, so the name identifier of the Edit page is assigned as Baicizhan0_1.
Since the UI pages from real-world apps may change as the app evolves. During exploration, if the head node stored several days before is simply approximated to the current head node, UI changes may be mistaken for the edge’s action, causing data errors. Therefore, for each page, the algorithm will save screenshots and XML documents of each step in its entire “action trace”. Taking “Baicizhan0_1_25” in Figure 2 as an example, this node contains 3 UI pages, 3 XML documents, and an “action trace” consisting of 2 actions.
We adopt the breadth-first method (BFS) to explore apps. Compared with the depth-first method (DFS), BFS better covers app functions and shortens the action sequence when exploring new nodes. As shown in Figure 2, the algorithm will first explore “Baicizhan0_10” instead of “Baicizhan0_1_25”. The task-oriented exploration method Yang et al. (2023) heavily relies on the performance of VLM. However, current VLMs can be costly and may perform poorly with third-party apps. In addition, the task-oriented method may cause some infrequently used pages and app functions to be overlooked.
@@ 3.3 Method Optimization
To improve exploration efficiency, we propose a “unique page” mechanism. Every time a new page is explored, we use BM25 Robertson et al. (2009) to retrieve the top 5 nodes in the current app graph that are closest to the XML document of the page. The algorithm compares the new page with each of these five pages to determine if they are similar pages. The threshold of the similar coefficient is Element Diff<5 & Pixel Diff <30%. Here, Element Diff is the number of different elements between two UI pages and Pixel Diff is the pixel difference between two screenshots. If no similar page is found in the current graph, the new page will be treated as a unique page and added to the app graph. As shown in Figure 2, click the “Back” button on “Baicizhan0_1_24”, and the generated “Baicizhan0_1_24_113” and “Baicizhan0” are equivalent pages. We add a directed edge from the previous page “Baicizhan0_1_24” to the similar page “Baicizhan0” in the graph. The algorithm will not treat the “Baicizhan0_1_24_113” as a new explorable node.
The benefits of this mechanism are threefold: 1. This greatly reduces the exploration space of each app. Taking “ctrip” as an example, our exploration process produced 187,079 UI pages with an average steps of 6.5. Without the “unique page” mechanism, pages of this magnitude cannot even cover all possibilities of 4-step exploration. 2. This converts the tree structure exploration results into a graph structure. Different pages can reach “Baicizhan0_1” by clicking “Edit”, “OK”, and scrolling. This helps the agent learn the functions of different UI elements. 3. This helps prevent the occurrence of cyclic action sequences. The “unique page” mechanism can detect and prune them.
To balance the distribution of different actions in the dataset, during random walks, we give priority to the input action. We provide 10 related keywords for each app. When executing the input action, the algorithm can randomly select a keyword to input. For scroll actions, the algorithm can choose a direction to scroll from “up, down, left, and right”.
@@ 3.4 Dataset Statistics
Among the 49 selected apps, we ensure that each main category in AppStore contains at least 2 apps. Figure 4 shows the data distribution of the Mobile3M dataset. The most common application categories in the dataset are “Travel”, “Living” and “Shopping”. As shown in the figure, Mobile3M covers multiple categories, and the amount of data in each category is relatively balanced, which ensures that the dataset is versatile and diverse.
@@ 4.3 Model Architecture
Table 4: Main Result(%). Suboptimal results are marked with an underline. The BC results are from (Sun et al., 2022), results with * are from Ding (2024), and results with # stem from (Zhan and Zhang, 2023). Considering the testing costs, Qwen-VL-Max and GPT-4o were conducted on a random sample of 10% of Auto-UI.
@@ 5.6 Pre-training Results
5.7 Language Environment Analysis.
In the main experiments, we translate the instructions into Chinese to align with the language we use in the pre-training stage. However, Qwen-vl-chat’s navigation capabilities may be influenced when the environment changes back to English. Therefore, we supplement an experiment on the original Auto-UI dataset to compare the performance loss caused by the language environment.
Model General Install Single Google apps Web
MobileVLM_cn 68.87 79.49 67.38 74.85 70.85
MobileVLM_en 69.58 79.87 81.24 74.72 71.70
Table 7: Performance Comparison of Different Languages. CN means instructions are translated into Chinese, while EN means not.
From Table 7, it can be seen that most of the English results are very close to the Chinese results, except for some metrics. This shows that MobileVLM is a basic model for performing complex mobile end-to-end navigation tasks in a Chinese-English bilingual environment.
5.8 Resource Consumption Analysis
The inference consumption of MobileVLM under 4-bit quantization is approximately 23GB of GPU memory. Due to the need for memory to store intermediate computation results and .eval during inference, the memory requirement is generally 2-3 times that of the GPU memory. Therefore, if running MobileVLM on the device side (directly on a real phone to control apps), the model inference would require at least an 8-core CPU, 46GB of RAM, and 23GB of dedicated GPU memory.
We designed an additional experiment: over a period of 4 minutes, with each 30-second interval as a timestamp, we recorded the memory and GPU consumption during model inference. See the table 8 below:
Timestamp (s) RAM (MB) GPU Memory (MB)
30 2009.87 18502.0
60 3923.33 20310.0
90 3940.03 20798.0
120 18438.99 21550.0
150 20312.5 23128.0
180 28906.25 23128.0
210 39471.97 23128.0
240 38438.99 23128.0
Table 8: Memory Usage Over Time
6 Discussion
6.1 Can MobileVLM be applied in real applications?
There are three main challenges to addressing the practical deployment issues of the model: 1. Resource issues, 2. Inference speed, 3. Permission issues. Below, we will discuss them in detail:
  1. 1.
    Resource Issues: As we know, current on-device large models typically range from 700M to 4.5B parameters. For example, Apple’s OpenELM has 2.7B parameters, and Microsoft’s Phi-3-mini has 3.8B parameters. Both models are specifically designed for mobile terminals. Our model has 9.8B parameters, which presents a significant resource requirement gap for on-device deployment. However, the smaller parameter models mentioned also have performance demand gaps.
  2. 2.
    Inference Speed Issues: Real-world mobile usage scenarios require quick model responses. For instance, common human-computer interaction validation requires completing clicks or swipes within a specified time. Additionally, when I need to view a scrolling advertisement, I must click it just as the target ad page scrolls from the background to the main interface. Both scenarios demand that the model completes inference within the specified time.
  3. 3.
    Permission Issues: As introduced in the above sections, VLM primarily relies on Appium and ADB to control the mobile device. When the model needs to be directly deployed on the phone to control apps, the dependent app, such as Siri, needs to have at least system-level permissions. This means that the app must have the current system’s system signature. However, for most closed-source mobile operating systems, granting system-level signatures to third-party apps is almost impossible.
@@ 7 Conclusion
Acknowledgements
We thank the Xiaomi SmartPhone Department of Xiaomi Technology Corporation for their emulator support for this project. This work was supported by the NSFC (U2001212, 62032001, and 61932004).
@@ A.3 Few-shot Prompt
For the Grounding Task, we use this few-shot prompt to guide the VLMs in answering the questions: {mdframed}[linewidth=1pt,linecolor=black] Here are three examples and a question. You need to help me find the location of the text I’m looking for in the image and output its bounding box. Please note that the coordinates are relative to the top-left corner of the image. Here are the three examples:
@@
For the Navigation Task, we use this few-shot prompt to guide the VLMs in answering the questions:
{mdframed}
[linewidth=1pt,linecolor=black] Here are two examples and a question. You need to tell me which control to interact with to navigate from the first image to the second image. There are three actions: click, input, and scroll. Here are the two examples:
@@
For the OCR Task, we use this few-shot prompt to guide the LLMs in answering the questions: {mdframed}[linewidth=1pt,linecolor=black]
@@
    Question: Image: <img>/home/corpus/test_515/few_shot/ocr/ctrip0_1_36_313_3617_1988_8 364_4440_1566/ctrip0_1_36_313_3617_1988_836
    4_4440_1566-screen.png</img> Explain how the elements on the current page support its content and functionality.
@@
    Question: Image: <img>/home/corpus/test_515/few_shot/ocr/baicizhan0_1_24_113_159_165
    _156_206/baicizhan0_1_24_113_159_165_
    156_206-screen.png</img> Summarize the layout and interaction methods of each part of the page.
@@
{mdframed}
[linewidth=1pt,linecolor=black] <ref>Added</ref><box>[613,524][682,558]</box>
@@ B.1 Background
B.2 MAU of selected APPs
To demonstrate that the selected apps, although few in number, cover the majority of daily use cases for Chinese users, we can use two key metrics: download volume and Monthly Active Users (MAU). For example, we will use data from the travel and music categories collected from the Tencent App Store.
Table 9: Travel Applications Statistics
Application Downloads (millions) Ratio
Ctrip 38000 48.17%
Qunar 35000 44.36%
Tongcheng 2600 3.30%
Fantawild 165.5 0.21%
Fliggy 8008 10.15%
Mafengwo 2318 2.93%
Aowei 95 0.12%
Tuniu 8082 10.25%
CYTS 62.3 0.08%
Disney 282.2 0.36%
Spring Tour 78.6 0.10%
Table 10: Music Applications Statistics
Application Downloads (millions) Ratio
QQ Music 274000 35.24%
Kugou 292000 37.56%
Qishui 257.9 0.03%
Kuwo 122000 15.69%
Tomato Changting 142.4 0.02%
Huisen 43 0.01%
Bodian 211.1 0.03%
Migu 27000 3.47%
Qianqian Queting 7872.2 1.01%
Apple Music 23.9 0.01%
Free 6.6 0.01%
From the table 10 and 10, we can see that although the number of apps we selected is relatively small, they account for over 70% of the usage. Additionally, Mobile3M includes a wide range of app categories that are sufficient to cover users’ daily life scenarios. The table below shows the number of active users in October 2023 in millions. The bolded sections are the apps we have chosen. The Monthly Active Users (MAU) data further supports the representativeness of these selected apps.
Table 11: Travel Applications Statistics
Application MAU (millions) Ratio
Ctrip 6964.7 51.96%
Qunar 2761 20.60%
Fliggy 1868.3 13.94%
Tongcheng 574 4.28%
Mafengwo 459 3.42%
Huazhu Club 351 2.62%
Zhixing 120 0.90%
Booking 85 0.63%
Ctrip Business 116 0.87%
State Grid Business 104 0.78%
Table 12: Music Applications Statistics
Application MAU (millions) Ratio
QQ Music 1869.7 32.09%
Kugou 1850.1 31.75%
Kuwo 929 15.94%
Qishui 327.1 5.61%
Himalaya 680.7 11.68%
Qingting 60.5 1.04%
Maoer 51.9 0.89%
Xiaoyuzhou 39 0.67%
Lizhi 19.8 0.34%
B.3 Graph Structure
@@
B.4 Category Details
Figure 7 shows a two-layered ring chart. The inner ring represents the data proportion by category, while the outer ring indicates the specific quantity proportion of each app. From this figure, it can be seen that, except for a few apps, most categories have a uniform distribution. For detailed quantity statistics, refer to Figure 12.
Figure 12: The data distribution in Mobile3M.